# A novel efficient structure-preserving exponential integrator for Hamiltonian systems

### v1

@@ 2 The new linearly implicit structure-preserving symmetric method
For the semilinear system (2), any exponential integrator with the form
@@
where $A=QM$, and ^$\hat{\nabla}U(x_{n},x_{n+1},\cdots,x_{n+k})$ is a discretiztion of the true gradient $\nabla U(x_{n})$, satisfies the following equation
@@
The proof of this corollary follows similarly as the proof of Lemma 3.
@@
For a homogeneous function $U(x)$, EKahan scheme (16) nearly preserves the exact energy and has the following error estimation:
$+\bar{U}(x_{n},x_{n+1},\dots,x_{n+k-2},x_{n+k-1},x_{n},x_{n+k}-x_{n+k-1})$
$-2\bar{U}(x_{n},x_{n+1},\dots,x_{n+k-1},x_{n+k},\frac{x_{n+k}-x_{n}}{k}),$
where $H_{n}=\frac{1}{2k}\sum\limits_{i=0}^{k-1}x_{n+i}^{\mathsf{T}}Mx_{n+i}+\bar{U}(x_{n},x_{n+1},\dots,x_{n+k-2},x_{n+k-1},x_{n},x_{n+k-1})$ is the discrete energy at $t=t_{n}$, and $\bar{U}(\cdot,\cdot,\dots,\cdot,\cdot)$ is a symmetric $k+2$ linear function defined via the $k+2$ tensor associated with the homogeneous function $U(x)$ and satisfies $\bar{U}(x,x,\dots,x,x)=U(x)$.
The proof of this corollary proceeds analogously to that of Theorem 2. For polynomial nonlinear functions $U(x)$ of order greater than three, the energy error can no longer be expressed in the simple form presented in Theorem 2; nevertheless, Corollary 3 demonstrates that the energy error over a single time step can be effectively controlled by the corresponding change in the numerical solution, thereby ensuring the boundedness of the energy error.
@@
For a nonhomogeneous polynomial $U(x)$, we can extend it to a homogeneous function by introducing an auxiliary variable $x^{0}$ to the original vector $x$. Denote the extended variable as $\bar{x}=(x^{0},x^{1},\cdots,x^{m})^{\mathsf{T}}$, where $x^{i}$, $i=1,\cdots,m$, are the components of the vector $x$. The nonhomogeneous cubic polynomial $U(x)$ is extended to the homogeneous function $\bar{U}(x^{0},x^{1},\cdots,x^{m})$ such that $\bar{U}(1,x^{1},\cdots,x^{m})=U(x)$. Furthermore, to preserve the structural properties of the original system in the extended formulation, we also augment the skew-symmetric matrix $Q$ to $\bar{Q}$ and the symmetric matrix $M$ to $\bar{M}$ by introducing an extra row and column of zero elements. Notably, $\bar{Q}$ remains skew-symmetric. Furthermore, applying the EKahan method to the modified differential equation
@@
produces the same update for $x$ as applying the EKahan method directly to the original reduced system (2). Therefore, all results previously established for systems with homogeneous polynomial nonlinearities extend naturally to those involving nonhomogeneous polynomial terms.
@@ 3 Numerical test
We define the global error by
$\mathcal{E}_{G}=\max\limits_{n\geq 0}\|y_{n}-y(t_{n})\|,$
and the energy error at time $t_{n}=t_{0}+nh$ by
$\mathcal{E}_{H}(n)=\lvert H_{n}-H(y(t_{n}))\rvert,$
where $y(t_{n})$ is the reference exact solution and $H_{n}$ is the discrete energy defined in Theorem 2. The reference exact solution is obtained using the 6-order continuous Runge-Kutta (CRK) method [31] with the form
@@
and the integrals are evaluated exactly by the 5-point GL quadrature. For both LIEEP method and EKahan scheme (16), multiple initial starting points are required. These starting values are obtained using the aforementioned reference exact solution. All numerical experiments were conducted using MATLAB (R2022b) on a 2020 MacBook Air equipped with an Apple M1 processor and 16 GB of unified memory.
Henon-Heiles system
The Henon-Heiles model [32] has been used to describe stellar motion in a long time course of gravitational potential energy tracking of galaxies with columnar symmetry. It has the following form
$\left\{\begin{array}[]{ll}{\ddot{q}_{1}=-q_{1}-2Dq_{1}q_{2}},\\ {\ddot{q}_{2}=-q_{2}-Dq_{1}^{2}+Cq_{2}^{2}},\\ \end{array}\right.$
@@
$\dot{y}=Q\nabla H,\quad H(y)=\frac{1}{2}y^{\mathsf{T}}My+U(y),$ (17)
where $y=(q_{1},q_{2},p_{1},p_{2})^{\mathsf{T}}$, $Q$ is the canonical skew-symmetric matrix, $M=I_{4}$ denotes the $4\times 4$ identity matrix, and the potential function is defined as $U(y)=q_{1}^{2}q_{2}-\frac{1}{3}q_{2}^{3}$.
Applying the EKahan scheme to equation (17) gives
$y_{n+1}=e^{hQM}y_{n}-h\phi(hQM)Q[\frac{1}{2}\nabla{U}(y_{n})-2\nabla U(\frac{y_{n+1}+y_{n}}{2})+\frac{1}{2}\nabla U(y_{n+1})],$
and we can get the accurate form of the exponential function
@@
Consider the initial data $y_{0}=(0,-0.082,0,0)^{\mathsf{T}}$, $T=100$ and time step size $h_{i}=0.02/2^{i},\;i=0,1,\dots,4$. Figure 1(a) demonstrates that the EAVF method preserves the discrete energy exactly and the energy errors of all linearly implicit methods remain bounded, varying from $10^{-18}$ to $10^{-7}$. Figure 1(b) verifies the result about the energy deviation at each time step by EKahan method shown in Theorem 2.
Figure 1: Energy behaviour for the Henon-Heiles equation with $T=100$ and $h_{0}=0.02$. Left: energy error computed by all schemes; right: the error of the true and deduced energy deviation by EKahan method shown in Theorem 2.
@@
Figure 2: Global error and computational time of different methods for the Henon-Heiles equation with $T=100$ and different time step sizes.
@@
where the independent variables $(x,t)\in[0,L]\times[0,T]$, the parameter $\epsilon>0$, the internal damping coefficient $\beta\geq 0$, and the external damping coefficient $\gamma\geq 0$. By introducing $\frac{\partial u}{\partial t}=v$, we can reformulate the second-order partial differential equation into the following first-order differential equations
$\left\{\begin{array}[]{ll}{\partial_{t}u=v,}\\ {\partial_{t}v=\beta\frac{\partial^{2}v}{\partial x^{2}}+\frac{\partial^{2}u}{\partial x^{2}}(1+\epsilon(\frac{\partial u}{\partial x})^{p})-\gamma v-m^{2}u}.\\ \end{array}\right.$
Setting $y=[u^{\mathsf{T}},v^{\mathsf{T}}]^{\mathsf{T}}$, we can rewrite the above equation into the following Hamiltonian form
@@
$Q=\left[\begin{array}[]{cc}0&I\\ -I&\beta D-\gamma I\\ \end{array}\right],M=\left[\begin{array}[]{cc}m^{2}I-D&0\\ 0&I\\ \end{array}\right],U(y)=\sum_{j=0}^{N-1}\frac{\epsilon}{6}(\frac{u_{j+1}-u_{j}}{\Delta x})^{3}.$
@@
$u_{j}(0)=q_{j}(0),\quad v_{j}(0)=\dot{q}_{j}(0),$
where the function $q_{j}(t)$ takes the form
$q_{j}(t)=5\sum_{k\in\{32,96\}}\ln\frac{1+e^{2(\alpha(j-k)+t\sinh\alpha)}}{1+e^{2(\alpha(j-k-1)+t\sinh\alpha)}}$
@@
We consider three different settings of damping coefficients: i) a conservative system with $\gamma=0$ and $\beta=0$, ii) a dissipative system with $\gamma=0.1$ and $\beta=0$, and iii) a dissipative system with $\gamma=0$ and $\beta=2$. Figure 3 shows that EKahan method preserves the energy structure of the equation in both conservative and dissipative settings:
  1. i)
@@
  3. iii)
    the energy of all methods drop along time for systems with a positive internal or external damping coefficient in Figure 3(c) and 3(d).
@@
Figure 3: Energy behaviour for the FPU system with $p=1$, $T=100$, and $h_{2}=0.25$. Figure 3(a): the energy error computed by all schemes for the conservative system; Figure 3(b): the error of the true and deduced energy deviation by EKahan method shown in Theorem 2 for the conservative system; Figure 3(c) and Figure 3(d): the energy computed by all schemes for the two dissipative systems.
We test the convergence order and computational efficiency for all methods with those three settings of damping coefficients and the results are consistent. We report the results for the conservative equations in Figure 4. Figure 4(a) shows that Ekahan method is second-order and and Figure 4(b) shows that EKahan method is more efficient than all other methods. In addition, we observe that Kahan method is less accurate than any exponential integrator although it is shown to be more computationally effective than EAVF method. In Figure 5, we illustrate the numerical solutions given by EKahan method. Figure 5(a) shows that EKahan method captures the conservative property of the solution, exhibiting coherent traveling wave structures that preserve their shape and amplitude over time. Figure 5(b) shows that EKahan method captures the decay of the wave profiles as energy dissipates internally. Figure 3(c) and 5(b) demonstrate that although EKahan is designed for conservative equations, it can also capture the dissipative nature for the dissipative systems.
@@
Figure 5: Numerical solutions of the FPU system computed over $t\in[0,500]$ with time step need confirmation $h_{2}=0.25$.
@@
We consider $p=2$, i.e., the potential $U(y)$ is a quartic homogeneous polynomial function. We employ identical initial boundary condition and computational parameters to the $p=1$ case, with the exception of setting $\epsilon=100$. Figure 6(a) illustrates that the EAVF method preserves the discrete energy exactly and the energy errors of all linearly implicit methods remain bounded and oscillated. Figure 6(b) verifies the formula of the energy deviation by EKahan method shown in Corollary 3.
Figure 6: Energy behaviour for the FPU system with $p=2$, $\beta=\gamma=0$, $T=100$ and $h_{2}=0.25$. Left: the energy error computed by all schemes; right: the error of the true and deduced energy deviation computed by EKahan method shown in Corollary 3.
@@
$u_{t}=-\partial_{x}\frac{\delta\mathcal{H}}{\delta u},\quad\frac{\delta\mathcal{H}}{\delta u}=\frac{1}{2}u^{2}+u_{xx}+u_{yy}.$
@@
$H(u)=\sum_{j=0}^{N_{y}}\sum_{i=0}^{N_{x}}\left(\frac{(\delta_{x}^{+}u_{i,j})^{2}+(\delta_{x}^{-}u_{i,j})^{2}}{4}+\frac{(\delta_{y}^{+}u_{i,j})^{2}+(\delta_{y}^{-}u_{i,j})^{2}}{4}+\frac{u_{i,j}^{3}}{6}\right)\Delta x\Delta y.$
Here, $\delta_{x}^{+}$ and $\delta_{x}^{-}$, similarly $\delta_{y}^{+}$ and $\delta_{y}^{-}$, denote the forward and backward finite difference operators in the $x$- and $y$-directions, respectively. First-order and second-order central difference operators are employed to approximate the first and second spatial derivatives, respectively. The corresponding discrete differential operators are denoted by $D_{x}^{(1)},D_{x}^{(2)},D_{y}^{(1)},$ and $D_{y}^{(2)}$. Using these notations, the semi-discrete system can be compactly written as
$\dot{U}=-D_{x}^{(1)}\left(\frac{1}{2}U.^{2}+D_{x}^{(2)}U+D_{y}^{(2)}U\right),$
@@
Numerical simulations are carried out on the spatial domain $[0,L]\times[0,L]$ with $L=6$ and the temporal domain $[0,T]$ with $T=8$, subject to periodic boundary conditions. The spatial discretization employs a uniform grid with $N=33$ points in each spatial direction, resulting in a mesh size of $\Delta x=\Delta y=L/(N-1)$. For time integration, different time steps $h_{i}=0.01/2^{i+1},\;i=1,2,3,4$ are adopted. We consider the initial condition
@@
Figure 8(a) demonstrates that the energy errors of all linearly implicit methods remain bounded and oscillated. Figure 8(b) verifies the result about the energy deviation at each time step by EKahan method shown in Theorem 2.
Figure 8: Energy behaviour for the Zakharov-Kuznetsov equation with $T=8$ and $h_{1}=0.0025$. Left: the energy error computed by EKahan, Kahan and LIEEP schemes; right: the error of the true and deduced energy deviation by EKahan shown in Theorem 2.
In Figure 9, we consider the global error and the computational cost using different step sizes $h_{i}$. Figure 9(a) demonstrates the second-order convergence of the EKahan method, and Figure 9(b) reveals its superior computational efficiency compared to Kahan method and LIEEP method.
@@
Figure 10 illustrates the solution of the Zakharov-Kuznetsov equation. In Figure 10(a), we observe that the initial condition exhibits a regular, periodic pattern composed of multiple interacting modes. As time progresses to $t=8$, the solution evolves into a more localized, bell-shaped structure, reflecting the nonlinear dispersive nature of the equation; as shown in Figure 10(b). Despite the significant structural changes in the solution profile, the total energy of the system remains nearly constant throughout the simulation, demonstrating the excellent long-time stability.

### v2

@@ 2.2 The multistep EKahan scheme for higher-degree polynomial Hamiltonians
For the semilinear system (2), any exponential integrator of the form
@@
where $\hat{\nabla}U(x_{n},x_{n+1},\cdots,x_{n+k})$ is a discretization of the true gradient $\nabla U(x_{n})$, satisfies
@@
For a homogeneous function $U(x)$ of degree $k+2$, EKahan scheme (19) satisfies the stepwise identity
$\begin{split}H_{n+1}-H_{n}=&\bar{U}(x_{n+1},x_{n+2},\dots,x_{n+k-1},x_{n+k},x_{n+k},x_{n+1}-x_{n})\\ &+\bar{U}(x_{n},x_{n+1},\dots,x_{n+k-2},x_{n+k-1},x_{n},x_{n+k}-x_{n+k-1})\\ &-2\bar{U}\Bigl(x_{n},x_{n+1},\dots,x_{n+k-1},x_{n+k},\frac{x_{n+k}-x_{n}}{k}\Bigr),\end{split}$ (29)
where
$H_{n}=\frac{1}{2k}\sum_{i=0}^{k-1}x_{n+i}^{\mathsf{T}}Mx_{n+i}+\bar{U}(x_{n},x_{n+1},\dots,x_{n+k-2},x_{n+k-1},x_{n},x_{n+k-1})$
is the discrete energy at $t=t_{n}$, and $\bar{U}(\cdot,\dots,\cdot)$ is a symmetric $(k+2)$-linear form associated with $U$ satisfying $\bar{U}(x,\dots,x)=U(x)$.
@@
For a nonhomogeneous polynomial $U(x)$, we homogenize it by introducing an auxiliary variable $x^{0}$ and an extended state $\bar{x}=(x^{0},x^{1},\ldots,x^{m})^{\mathsf{T}}$ so that $\bar{U}(1,x)=U(x)$. To preserve the structural properties of the original system, we also augment $Q$ to $\bar{Q}$ and $M$ to $\bar{M}$ by adding a zero row and column. Notably, $\bar{Q}$ remains skew-symmetric and $\bar{M}$ remains symmetric. Furthermore, applying the EKahan method to
@@
produces the same update for $x$ as applying the EKahan method directly to the original system (2). Therefore, all results previously established for systems with homogeneous polynomial Hamiltonians extend naturally to those involving nonhomogeneous polynomials.
@@ 3 Numerical test
We define the global error at time $t_{n}=t_{0}+nh$ as
$\mathcal{E}_{G}=\max_{n\geq 0}\|y_{n}-y(t_{n})\|,$
where $y(t_{n})$ is the reference exact solution. For EKahan, we measure the residual of the stepwise energy-increment identity by
$e_{n}=H_{n+1}-H_{n}-G_{n},$
where $G_{n}=U(x_{n+1}-x_{n})$ for cubic Hamiltonians and $H_{n}$ is the discrete energy defined in Theorem 2. For higher-degree polynomial Hamiltonians, $G_{n}$ is given by the right-hand side of equation (29) and $H_{n}$ is defined as in Corollary 3. The energy error is defined as
$\mathcal{E}_{H}(n)=\lvert H_{n}-H(y(t_{n}))\rvert$. (30)
The reference exact solution is obtained using the 6-order continuous Runge-Kutta (CRK) method [14] with the form
@@
and the integrals are evaluated exactly by the 5-point GL quadrature. For both LIEEP method and EKahan scheme (19), multiple initial starting points are required. These starting values are obtained using the aforementioned reference exact solution. The matrix exponential functions $e^{khQM}$ and $\phi(khQM)$ are computed using the MATLAB package presented in [2] when no explicit form is given, where Padé approximations are employed. We solve the linear systems arising from each step of the linearly implicit schemes (EKahan, Kahan and LIEEP) by MATLAB’s backslash operator. The nonlinear updates required for the reference solutions and the EAVF numerical solutions are computed via a fixed-point iteration, terminated when the 2-norm of the difference between successive iterates falls below a tolerance of $10^{-14}$. We also tested tolerances from $10^{-14}$ to $10^{-6}$ for EAVF and observed that EKahan is consistently the most efficient method across our test problems, except for the FPU system with $p=2$, where EKahan is not as efficient as EAVF when a bigger tolerance is considered. All numerical experiments were conducted using MATLAB (R2022b) on a 2020 MacBook Air equipped with an Apple M1 processor and 16 GB of unified memory.
Hénon-Heiles system
The Hénon-Heiles model [11] has been used to describe stellar motion in a long time course of gravitational potential energy tracking of galaxies with columnar symmetry. It has the following form
[…]
@@
$\dot{y}=Q\nabla H,\quad H(y)=\frac{1}{2}y^{\mathsf{T}}My+U(y),$
where $y=(q_{1},q_{2},p_{1},p_{2})^{\mathsf{T}}$, $Q$ is the canonical skew-symmetric matrix, $M=I_{4}$ denotes the $4\times 4$ identity matrix, and the potential function is defined as $U(y)=q_{1}^{2}q_{2}-\frac{1}{3}q_{2}^{3}$. The exponential function can be computed explicitly by
@@
Consider the initial data $y_{0}=(0,-0.082,0,0)^{\mathsf{T}}$, $T=100$ and time step size $h_{i}=0.02/2^{i},\;i=0,1,\dots,4$. Figure 1(a) demonstrates that the EAVF method preserves the discrete energy exactly and the energy errors of all linearly implicit methods remain bounded, varying from $10^{-18}$ to $10^{-7}$. Figure 1(b) verifies the result about the energy deviation at each time step by EKahan shown in Theorem 2. Furthermore, we verified numerically that the per-step energy increment of EKahan scales as $O(h^{3})$, consistent with the stepwise identity in Theorem 2, and we omitted the corresponding figure due to space limitations.
Figure 1: Hénon-Heiles equation with $T=100$ and $h_{0}=0.02$. Left: energy error defined in (30); right: the residual of equation (16) in Theorem 2.
@@
Figure 2: Global error and computational time of different methods for the Hénon-Heiles equation with $T=100$ and different time step sizes.
@@
where the independent variables $(x,t)\in[0,L]\times[0,T]$, the parameter $\epsilon>0$, the internal damping coefficient $\beta\geq 0$, and the external damping coefficient $\gamma\geq 0$. By introducing $\frac{\partial u}{\partial t}=v$ and setting $y=[u^{\mathsf{T}},v^{\mathsf{T}}]^{\mathsf{T}}$, we can rewrite the above equation into the following Hamiltonian form
@@
$Q=\left[\begin{array}[]{cc}0&I\\ -I&\beta D-\gamma I\\ \end{array}\right],M=\left[\begin{array}[]{cc}m^{2}I-D&0\\ 0&I\\ \end{array}\right],U(y)=\sum_{j=0}^{N-1}\frac{\epsilon}{6}(\frac{u^{j+1}-u^{j}}{\Delta x})^{3}.$
@@
$u^{j}(0)=q^{j}(0),\quad v^{j}(0)=\dot{q}^{j}(0),$
where the function $q^{j}(t)$ takes the form
$q^{j}(t)=5\sum_{k\in\{32,96\}}\ln\frac{1+e^{2(\alpha(j-k)+t\sinh\alpha)}}{1+e^{2(\alpha(j-k-1)+t\sinh\alpha)}}$
@@
We consider three different settings of damping coefficients: 1) a conservative system with $\gamma=0$ and $\beta=0$, 2) a dissipative system with $\gamma=0.1$ and $\beta=0$, and 3) a dissipative system with $\gamma=0$ and $\beta=2$. We emphasize that our theoretical results are derived for the conservative Hamiltonian case, where the stepwise energy-increment identity in Theorem 2 applies. The two dissipative settings are included as additional numerical tests to assess whether EKahan captures the expected energy-decay behavior when damping is present. Figure 3 summarizes the observed energy behavior:
@@
  • in the damped cases ($\gamma>0$ or $\beta>0$), the computed energy decreases over time, consistent with the dissipative nature of the underlying dynamics; see Figures 3(c) and 3(d).
@@
Figure 3: FPU system with $p=1$, $T=100$, and $h_{2}=0.25$. Figure 3(a): the energy error defined in (30) computed by all schemes for the conservative system; Figure 3(b): the residual of equation (16) in Theorem 2 for the conservative system; Figure 3(c) and Figure 3(d): the energy defined in Theorem 2 of all schemes for the two dissipative systems.
We test the convergence order and computational efficiency for all methods with those three settings of damping coefficients and the results are consistent. We report the results for the conservative equations in Figure 4. Figure 4(a) shows that EKahan is second-order and and Figure 4(b) shows that EKahan is more efficient than all other methods. In Figure 5, we illustrate the numerical solutions given by EKahan method. Figure 5(a) shows that EKahan captures the conservative property of the solution, exhibiting coherent traveling wave structures that preserve their shape and amplitude over time. Figure 5(b) shows that EKahan captures the decay of the wave profiles as energy dissipates internally. Figure 3(c) and 5(b) demonstrate that although EKahan is designed for conservative equations, it can also capture the dissipative nature for the dissipative systems.
@@
Figure 5: Numerical solutions of the FPU system computed over $t\in[0,500]$ with time step $h_{2}=0.25$.
@@
When $p=2$, the potential function is a quartic homogeneous polynomial with the form
$U(y)=\sum_{j=0}^{N-1}\frac{\epsilon}{12}\left(\frac{u^{j+1}-u^{j}}{\Delta x}\right)^{4}.$
Define the spatial discrete difference by
$w^{j}=\frac{u^{j+1}-u^{j}}{\Delta x},\quad j=0,\ldots,N-1.$
Following equation (18), the polarized gradient is given by
$\left(\nabla_{\mathrm{K}}U(y_{n},y_{n+1},y_{n+2})\right)^{j}=\frac{\epsilon}{3\Delta x}\left(w^{j-1}_{n}w^{j-1}_{n+1}w^{j-1}_{n+2}-w^{j}_{n}w^{j}_{n+1}w^{j}_{n+2}\right)$
for $j=0,\ldots,N-1$. The EKahan scheme then takes the form
$y_{n+2}=e^{2hQM}y_{n}+2h\phi(2hQM)Q\nabla_{\mathrm{K}}U(y_{n},y_{n+1},y_{n+2}).$
We employ identical initial boundary condition and computational parameters to the $p=1$ case, with the exception of setting $\epsilon=100$. Figure 6(a) illustrates that the EAVF method preserves the discrete energy exactly and the energy errors of all linearly implicit methods remain bounded and oscillated. Figure 6(b) verifies the formula of the energy deviation by EKahan method shown in Corollary 3.
Figure 6: FPU system with $p=2$, $\beta=\gamma=0$, $T=100$ and $h_{2}=0.25$. Left: the energy error (30) of all schemes; right: the residual of equation (29) in Corollary 3 by EKahan.
@@
$u_{t}=-\partial_{x}\frac{\delta\mathcal{H}}{\delta u},\quad\frac{\delta\mathcal{H}}{\delta u}=\frac{1}{2}u^{2}+u_{xx}+u_{yy}.$ (31)
@@
$H(u)=\sum_{j=0}^{N_{y}}\sum_{i=0}^{N_{x}}\left(\frac{(\delta_{x}^{+}u^{i,j})^{2}+(\delta_{x}^{-}u^{i,j})^{2}}{4}+\frac{(\delta_{y}^{+}u^{i,j})^{2}+(\delta_{y}^{-}u^{i,j})^{2}}{4}+\frac{(u^{i,j})^{3}}{6}\right)\Delta x\Delta y.$
Here, $\delta_{x}^{+}$ and $\delta_{x}^{-}$, similarly $\delta_{y}^{+}$ and $\delta_{y}^{-}$, denote the forward and backward finite difference operators in the $x$- and $y$-directions, respectively. First-order and second-order central difference operators are employed to approximate the first and second spatial derivatives in (31), respectively, and the corresponding matrices are denoted by $D_{x}^{(1)},D_{x}^{(2)},D_{y}^{(1)},$ and $D_{y}^{(2)}$. Using these notations, the semi-discrete system can be compactly written as
$\dot{U}=-D_{x}^{(1)}\left(\frac{1}{2}U^{2}+D_{x}^{(2)}U+D_{y}^{(2)}U\right),$
@@
Numerical simulations are carried out on the spatial domain $0,L]\times[0,L]$ with $L=6$ and the temporal domain $[0,T]$ with $T=8$, subject to periodic boundary conditions. The spatial discretization employs a uniform grid with $N=33$ points in each spatial direction, resulting in a mesh size of $\Delta x=\Delta y=L/(N-1)$. We consider different time steps $h_{i}=0.01/2^{i+1},\;i=1,\cdots,4$, and the initial condition
@@
Figure 8(a) demonstrates that the energy errors of all linearly implicit methods remain bounded and oscillated. Figure 8(b) confirms the stepwise energy-increment identity for EKahan established in Theorem 2.
Figure 8: Zakharov-Kuznetsov equation with $T=8$ and $h_{1}=0.0025$. Left: the energy error defined in (30); right: the residual of equation (16) in Theorem 2.
@@
In Figure 9, we consider the global error and the computational cost using different step sizes $h_{i}$. Figure 9(a) demonstrates the second-order convergence of the EKahan method, and Figure 9(b) reveals its superior computational efficiency compared to the other methods. Figure 10 illustrates the solution of the Zakharov-Kuznetsov equation. In Figure 10(a), we observe that the initial condition exhibits a regular, periodic pattern composed of multiple interacting modes. As time progresses to $t=8$, the solution evolves into a more localized, bell-shaped structure, reflecting the nonlinear dispersive nature of the equation; as shown in Figure 10(b). Despite the significant structural changes in the solution profile, the total energy of the system remains nearly constant throughout the simulation, demonstrating the excellent long-time stability.